# Learning Adaptive Dexterous Grasping from Single Demonstrations

### v1

@@ II-A Dexterous Manipulation
Dexterous manipulation is highly challenging due to the high degree of freedom of the end effector. Traditional methods that use trajectory optimization [2, 3, 4] require precise modeling of robot dynamics, which is often not feasible in the open world. In contrast, learning-based methods seek to obtain an observation-to-action mapping from data. Reinforcement learning methods [20, 21, 22] often require extensive exploration in simulation and specific reward engineering. In addition, the learned grasping policies from reward engineering could be unnatural when compared with human. Imitation learning methods [23, 24] are able to leverage real-world data collected with teleoperation systems [6, 7, 8], but real-robot data is costly to acquire and the policy is limited by the quality and amount of expert demonstrations. In this work, we aim to leverage a single real-world human demonstration to accelerate reinforcement learning, alleviate the burden for reward engineering, and acquire human-like grasp poses while using a learning curriculum to enable generalization beyond the demonstrated object configuration.
@@ II-B Learning from Human-Object Interaction Data
Human-object interaction datasets labelled with 3D human pose and 6-DoF object pose (e.g., [25, 26, 27]) are a rich repository for learning dexterous manipulation skills. [28] extracts robot and object poses from human videos for imitation learning but suffers from noisy state-action pairs, requiring hundreds of demonstrations and extensive reward engineering. To address these limitations, [10] introduces a trajectory mapping reward to learn from a single demonstration, leveraging trajectory augmentation and policy distillation for better generalization. However, its trajectory mapping reward is suboptimal as it enforces alignment with potentially imperfect human demonstrations. Other approaches learn to generate human-object interaction trajectories as reference trajectories in RL [29] or learn an imperfect policy from retargeted human demonstrations, refining it via RL or imitation learning [30]. Our work differs by learning a library of grasping skills using a trajectory following reward and a learning curriculum, enabling sample-efficient acquisition of human-like grasps and adaptive skill selection.
@@ II-C Language-Grounded Manipulation
An alternative approach to incorporating human preferences leverages Vision-Language Models (VLMs) alongside structured skill libraries [18, 38, 17, 19]. These methods use VLMs to interpret language instructions, extract preference-related information, and retrieve suitable skills from a predefined set. While effective in various manipulation tasks, their application to dexterous manipulation remains limited, particularly in adapting grasp poses based on user intent. This work extends this paradigm by constructing a skill library from human demonstrations and integrating a VLM to enable preference-driven grasp selection, bridging the gap between language understanding and dexterous control.
@@ III-A Overview
The environment then transitions to a new state $s_{t+1}$ according to the dynamics model $P$, and the robot receives a reward $r_{t}$. Each skill policy $\pi_{i}\in\Pi$ is optimized to maximize the expected cumulative reward given a reference trajectory extracted from a human demonstration $\tau_{i}$:
@@
where $H$ is the task horizon, and $\tau_{i}$ is a reference trajectory, which could be a video of a human grasping an object. The reference trajectory serves two purposes: (i) guiding reinforcement learning to improve sample efficiency, and (ii) encouraging the learned grasping strategy to align with human behavior.
@@ III-B Human Videos as Robot Demonstrations
To obtain human demonstrations, we use the DexYCB dataset, which provides real-world video recordings along with ground-truth hand and object poses for various objects, such as a sugar box, mustard bottle, and bleach cleanser [25]. For each object, the dataset includes demonstrations with at least two initial object positions and two distinct grasp poses. We seek to translate human demonstrations to robot trajectories to guide downstream reinforcement learning.
@@ III-C Reward Function Designs
Trajectory following reward. The reward guides the robot hand toward the pre-grasp stage. Existing work [10] enforces state-to-state alignment between the robot trajectory and the reference trajectory, which can be overly restrictive, as human demonstrations may contain suboptimal or infeasible motions. Instead, the agent should have the flexibility to reach key states using its own strategy while still leveraging human demonstrations. To achieve this, we propose a reward that measures the proximity of robot states and the reference trajectory, inspired by TR2 [47]. The reward at time step $t$ is defined as:
@@
[…]
@@ III-D Curriculum Learning
While the reward terms guide the learner to follow the reference trajectory and discover physically plausible robot trajectories to achieve the goal, the learned policy may struggle to generalize to varying initial object positions and orientations due to the diversity in demonstrations. To enhance generalization, we employ curriculum learning, progressively increasing task difficulty to expose the policy to a broader range of object configurations.
Initially, we use one fixed object’s pose for learning so that the policy quickly discovers feasible grasp poses. We then gradually randomize the initial pose, defined as a tuple of position and rotation $(p^{\text{pos}},p^{\text{rot}})$: $p^{\text{pos}}\sim\mathcal{U}(p_{\text{init}}^{\text{pos}}-\sigma P_{\text{max}},p_{\text{init}}^{\text{pos}}+\sigma P_{\text{max}}),\quad p^{\text{rot}}\sim\mathcal{U}(p^{\text{rot}}_{\text{init}}-\sigma\Theta_{\text{max}},p^{\text{rot}}_{\text{init}}+\sigma\Theta_{\text{max}})$ where $(p_{\text{init}}^{\text{pos}},p^{\text{rot}}_{\text{init}})$ is the original pose, $\mathcal{U}$ stands for uniform distribution, $P_{\text{max}}$ and $\Theta_{\text{max}}$ represent the maximum position displace in $x,y$ and the maximum rotation deviation around the $z$-axis, respectively. The noise factor $\sigma\in[0,1]$ controls the degree of variation, which is initialized as 00 and increased incrementally by $0.01$ until reaching $1$ after the performance exceeds a predefined success threshold $\zeta$. This ensures that the policy learns to generalize to diverse object configurations at a high sample efficiency.
@@ IV-A1 Evaluation Setup
Implementation. We use Proximal Policy Optimization (PPO) [52] for policy learning. For each object and one specific grasping pose, we first train the policy for 80M steps with a fixed initial object pose. After obtaining a stable policy, we continue training for an additional 120M steps with the proposed curriculum. Specifically, we apply randomization with an $xy$ translation range of ±$\pm 5$ cm and a rotation range of ±$\pm 30^{\circ}$ around the $z$-axis at the end of curriculum training. A success is rewarded when the object’s height is above $20$ cm. During testing, we evaluate the learned policies across different initial object configurations, using the same randomized strategy applied at the end of curriculum training.
@@ IV-A2 Main Results
We compare our method with the state-of-the-art, ViViDex [10], which uses a trajectory mapping reward to improve RL training, followed by imitation learning on rollouts for better generalization. As the official implementation is unavailable, we implement the reward and perform training for 200M steps. To ensure fairness, we use the same randomized initial object poses from our curriculum training for both training and testing. In addition, we implement a version of ViViDex with our proposed learning curriculum, which differs from AdaDexGrasp only in the reward function design, for a more controlled comparison.
@@
[…]
@@ IV-A4 Ablation Study of Reward Terms
We conduct ablation studies across eight grasping environments with fixed initial object poses to evaluate the impact of three reward terms. We compare our full pipeline with variations where each reward term is omitted. Without contact rewards, height rewards are given without contact checks. Without height rewards, we remove the height-proportional reward but retain the success reward when the object reaches the target height. The results are shown in Fig. II.
@@
[…]
@@ IV-A5 Effects of Curriculum Learning
We observe that a lower success threshold leads to a faster increase in noise level. However, when comparing the steps at which different settings achieve the same success rate after the noise level reaches 1, a higher success threshold results in greater training efficiency. This suggests that maintaining a high success rate while gradually introducing noise during curriculum learning yields better sample efficiency than rapidly increasing the noise level. Furthermore, direct training leads to a significantly lower success rate, underscoring the effectiveness of curriculum learning in improving the efficiency of RL training.
@@
[…]
Skill 1 $28$ $100$ 00 00 00
Skill 2 $72$ 00 $100$ 00 00
Skill 3 00 00 00 $100$ $100$
@@ IV-B2 Results
In terms of overall task execution success rate (Table IV), random selection often fails to select the appropriate skill. In contrast, VLM-based selection not only aligns well with human preferences but also achieves a high success rate in task execution.
@@ IV-C1 Evaluation Setup
Task setup. We validate the sim-to-real transfer of our policies on a real robot. We focus on three objects with varying geometry, appearance, and mass distributions: mustard-bottle, bleach-cleanser, and cracker-box. For each object, we use the first grasp pose and apply a random ±$\pm 20^{\circ}$ rotation and a ±$\pm 3$ cm position shift, repeating the experiment ten times.
@@
[…]
ViViDex[10] 00 00 00
@@ IV-C2 Results
As shown in Table V, our method consistently achieves strong performance, maintaining at least a $90\%$ success rate across objects. In contrast, the baseline fails on all three objects, which is expected given its near-zero success rate in simulation. Figure 5 presents qualitative results, illustrating the robot grasping objects with different poses.
V Conclusion
We introduced AdaDexGrasp, a framework for efficiently learning and adaptively selecting dexterous grasping skills. By combining human demonstrations, reinforcement learning with a trajectory following reward, and curriculum learning, AdaDexGrasp enables sample-efficient acquisition of diverse, human-like grasps. Additionally, a vision-language model allows for intuitive grasp selection based on task context. We demonstrate strong performance in both simulation and real-world experiments, paving the way for more adaptable and sample-efficient dexterous manipulation.

### v2

@@ II-A Dexterous Manipulation
Dexterous manipulation is highly challenging due to the high degree of freedom of the end effector. Traditional methods that use trajectory optimization [2, 3, 4, 20] require precise modeling of robot dynamics, which is difficult in the open world. In contrast, learning-based methods seek to obtain an observation-to-action mapping from data. Reinforcement learning methods [21, 22, 23, 24] often require extensive exploration in simulation and specific reward engineering. In addition, the learned grasping policies from reward engineering could be unnatural when compared with human. Imitation learning methods [25, 26] are able to leverage real-world data collected with teleoperation systems [6, 7, 8], but real-robot data is costly to acquire and the policy is limited by the quality and amount of expert demonstrations. In this work, we aim to leverage a single real-world human demonstration to accelerate reinforcement learning, alleviate the burden for reward engineering, and acquire human-like grasp poses while using a learning curriculum to enable generalization beyond the demonstrated object configuration.
@@ II-B Learning from Human-Object Interaction Data
Human-object interaction datasets with 3D human pose and 6-DoF object pose annotations (e.g., [27, 28, 29]) provide rich resources for learning dexterous manipulation. [30] extracts robot and object poses from human videos for imitation learning, but suffers from noisy state-action pairs and relies on hundreds of demonstrations and extensive reward tuning. To overcome these issues, [10] proposes a trajectory mapping reward for one-shot learning, using trajectory augmentation and policy distillation to improve generalization. However, this reward is suboptimal as it enforces alignment with potentially flawed demonstrations. Other methods generate human-object trajectories as RL references [31], or refine imperfect policies learned from retargeted demonstrations via RL or imitation [32]. In contrast, we learn a grasping skill library using a trajectory following reward and curriculum learning, enabling sample-efficient, human-like grasp acquisition and adaptive skill selection.
@@ II-C Language-Grounded Manipulation
An alternative approach to incorporating human preferences leverages Vision-Language Models (VLMs) [40, 41] with structured skill libraries [18, 42, 19, 17, 43]. These methods use VLMs to interpret language instructions, extract preference-related information, and retrieve suitable skills from a predefined set. While effective in various manipulation tasks, their application to dexterous manipulation remains limited, particularly in adapting grasp poses to user intent. This work extends the paradigm by building a skill library from human demonstrations and integrating a VLM to enable preference-driven grasp selection, bridging the gap between language understanding and dexterous control.
@@ III-A Overview
The environment then transitss to a new state $s_{t+1}$ according to the dynamics model $P$, and the robot receives a reward $r_{t}$. Each skill policy $\pi_{i}\in\Pi$ is optimized to maximize the expected cumulative reward given a reference trajectory extracted from a human demonstration $\tau_{i}$:
@@
where $H$ is the task horizon. The reference $\tau_{i}$ (e.g., a human grasping video) both guides learning for sample efficiency and promotes human-like behavior.
@@ III-B Human Videos as Robot Demonstrations
To obtain human demonstrations, we use the DexYCB dataset, which provides real-world video recordings along with ground-truth hand and object poses for various objects[27]. Each object has demonstrations with at least two initial positions and two grasp poses. We seek to translate human demonstrations to robot trajectories to guide downstream reinforcement learning.
@@ III-C Reward Function Designs
Trajectory following reward. The reward guides the robot hand toward the pre-grasp stage. Existing work [10] enforces state-to-state alignment between the robot trajectory and the reference trajectory, which can be overly restrictive, as human demonstrations may contain suboptimal or infeasible motions. Instead, agents should flexibly reach key states using their own strategies while leveraging demonstrations. To achieve this, we propose a reward that measures the proximity of robot states and the reference trajectory, inspired by TR2 [52]. The reward at time step $t$ is defined as:
@@
Here, state $s^{i,*},d^{i,*}$ denotes the relative pose of the $i$-th fingertip to the object, with $\text{pos},\text{rot}$ indicating position and rotation, respectively. $r_{\text{dist}}(s,d)$ is defined as: $r_{\text{dist}}(s,d)=1-\tanh(\text{dist}(s,d)),$ mapping the distance to a bounded reward that encourages the policy to reach states closer to the reference. In this paper, we refer to our reward and that in [10] as the trajectory following and trajectory mapping reward, respectively.
@@ III-D Curriculum Learning
While the reward terms guide the learner to follow the reference and discover feasible robot trajectories, the policy may struggle to generalize to diverse initial object poses due to variation in demonstrations. To improve generalization, we use curriculum learning, gradually increasing task difficulty to expose the policy to a wider range of object configurations.
Initially, we use one fixed object’s pose for learning so that the policy quickly discovers feasible grasp poses. We then gradually randomize the initial pose, defined as a tuple of position and rotation $(p^{\text{pos}},p^{\text{rot}})$: $p^{\text{pos}}\sim\mathcal{U}(p_{\text{init}}^{\text{pos}}-\sigma P_{\text{max}},p_{\text{init}}^{\text{pos}}+\sigma P_{\text{max}}),\quad p^{\text{rot}}\sim\mathcal{U}(p^{\text{rot}}_{\text{init}}-\sigma\Theta_{\text{max}},p^{\text{rot}}_{\text{init}}+\sigma\Theta_{\text{max}})$ where $(p_{\text{init}}^{\text{pos}},p^{\text{rot}}_{\text{init}})$ is the original pose, $\mathcal{U}$ stands for uniform distribution, $P_{\text{max}}$ and $\Theta_{\text{max}}$ represent the maximum position displace in $x,y$ and the maximum rotation deviation around the $z$-axis, respectively. The noise factor $\sigma\in[0,1]$ controls the degree of variation, which is initialized as 0 and increased incrementally by $0.01$ until reaching $1$ after the performance exceeds a predefined success threshold $\zeta$. This ensures that the policy learns to generalize to diverse object configurations at a high sample efficiency.
@@ IV-A1 Evaluation Setup
Implementation. We adopt Proximal Policy Optimization (PPO) [57] for policy learning. The robot observes the full simulation state, including all joint angles, and the poses of the hand, fingertips, and objects, represented in the robot arm base frame. The policy outputs joint angle deltas, indicating real-time changes for actuation. For each object and grasp pose, we train the policy for 80M steps with a fixed initial pose, followed by another 120M steps using the proposed curriculum. During curriculum training, we randomize object position within $\pm 5$ cm in $xy$ and rotation within $\pm 30^{\circ}$ around the $z$-axis. A success is defined as lifting the object above 20 cm. At test time, we evaluate the policy on varied initial object configurations, using the same randomization as in curriculum training.
@@ IV-A2 Main Results
We compare our method with the state-of-the-art ViViDex [10], which uses a trajectory mapping reward to aid RL training, followed by imitation learning on rollouts for better generalization. The observation and action spaces are the same as in AdaDexGrasp. For fair comparison, we trained both our reimplementation of ViViDex and AdaDexGrasp for 200M steps. Our proposed reward design differs in several aspects. In the pre-grasp stage, we reward object approach and contact. These are retained during manipulation, with an added reward for object motion alignment. To address sim-demo length mismatch, we use progress-based frame selection. The object-approach reward extends ViViDex with rotation-aware distance, while the contact reward follows Section III-C. To ensure fairness, we use the same randomized initial object poses from our curriculum training for both training and testing. In addition, we implement a version of ViViDex with our proposed learning curriculum, which differs from AdaDexGrasp only in the reward function design, for a more controlled comparison.
@@
SR ($\%$) $\uparrow$ mustard-bottle bleach-cleanser cracker-box sugar-box Avg
@@ IV-A4 Ablation Study of Reward Terms
We conduct ablation studies across eight grasping environments with fixed initial object poses. We compare our full pipeline with variations where each reward term is omitted. Without contact rewards, height rewards are given without contact checks. Without height rewards, we remove the height-proportional reward but retain the success reward when the object reaches the target height. The results are shown in Fig. II.
@@
SR ($\%$) $\uparrow$ mustard-bottle bleach-cleanser cracker-box sugar-box Avg
@@ IV-A5 Effects of Curriculum Learning
We observe that a lower success threshold causes a faster rise in noise level. However, when comparing the steps at which different settings reach the same success rate after the noise level hits 1, a higher threshold yields greater training efficiency. This suggests that maintaining a high success rate while gradually introducing noise during curriculum learning offers better sample efficiency than rapidly increasing noise. Furthermore, direct training results in a significantly lower success rate, underscoring the effectiveness of curriculum learning in improving RL training efficiency.
@@
P(skill $|$ task) ($\%$) T1 T2 T3 T4 T5
Skill 1 $28$ $100$ 0 0 0
Skill 2 $72$ 0 $100$ 0 0
Skill 3 0 0 0 $100$ $100$
@@ IV-B2 Results
In terms of overall task success rate (Table IV), random selection often fails to choose the appropriate skill. In contrast, VLM-based selection aligns with human preferences and achieves high task success.
@@ IV-C1 Evaluation Setup
Task setup. We validate the sim-to-real transfer of our policies on a real robot. We focus on three objects with varying geometry, appearance, and mass distributions: mustard-bottle, bleach-cleanser, and cracker-box. For each object, we use the first grasp pose and apply a random $\pm 20^{\circ}$ rotation and a $\pm 3$ cm position shift, repeating the experiment ten times.
@@
SR ($\%$) $\uparrow$ mustard-bottle bleach-cleanser cracker-box
ViViDex[10] 0 0 0
@@ IV-C2 Results
As shown in Table V, our method consistently achieves strong performance, maintaining at least a $90\%$ success rate across objects. In contrast, the baseline fails on all three objects, as expected given its near-zero success rate in simulation. Figure 5 shows qualitative results, illustrating the robot grasping objects with different poses.
V Conclusion & Limitations
We introduced AdaDexGrasp, a framework for efficiently learning and adaptively selecting dexterous grasping skills. By integrating human demonstrations, demonstration-guided reinforcement learning, curriculum learning, and VLM-based skill selection, AdaDexGrasp enables sample-efficient acquisition of diverse, human-like grasps. We demonstrate strong performance in both simulation and real-world experiments.
Our approach has a few limitations. First, as the skill library scales to include more grasp and object types, retrieving the appropriate policy becomes more susceptible to hallucinations by VLMs. Structured knowledge representations (e.g., symbolic) may help encode skill affordances [43] and support more reliable information update and reasoning [62]. Second, while we test the model on simple, ambiguous instructions, its robustness to more complex or noisy language inputs remains to be explored. These challenges open promising directions for future work.